# Under the null of valid specification, pre-tests of valid specification do not distort inference

### v1

@@ Abstract
Consider a parameter of interest, which can be consistently estimated under some conditions. Suppose also that we can at least partly test these conditions with specification tests. We consider the common practice of conducting inference on the parameter of interest conditional on not rejecting these tests. We show that if the tested conditions hold, conditional inference is valid but possibly conservative. This holds generally, without imposing any assumption on the asymptotic dependence between the estimator of the parameter of interest and the specification test.
@@ 1 Introduction
Theoretical guarantees on inference methods are usually given unconditionally. Yet, inference on a parameter of interest is rarely conducted in isolation of other statistical practices. In this paper, we consider the common practice of reporting confidence intervals or performing tests on a parameter of interest, $\beta_{0}$, only if specification tests are not rejected. For instance, one may proceed with a difference-in-difference (DID) estimator only if a pretrends test is not rejected. One may then wonder if pre-testing may distort inference. Let $\widehat{\beta}$ denote an estimator of $\beta_{0}$. This note shows that under the null that the conditions underlying the validity of $\widehat{\beta}$ hold, conditional inference is valid, albeit possibly conservative. This result holds under general conditions. In particular, no restriction on the asymptotic dependence between $\widehat{\beta}$ and the specification tests is needed. The main condition on these specification tests is that they are convex functionals. This holds for finite-dimensional $F$-tests but also for infinite-dimensional tests based on empirical processes such as (weighted) Kolmogorov-Smirnov or Crámer-von Mises tests. For finite-dimensional test statistics, our result is a direct application of the Gaussian correlation inequality (Royen, 2014). For infinite-dimensional ones, we first rely on an approximation before applying this inequality.
For instance, in a DID analysis, our result implies that under the null of parallel trends, inference conditional on not rejecting a pretrends test is at worst conservative, but can never be liberal, irrespective of the correlation structure between the DID and pre-trends estimators. This has implications for some recent discussions in that literature. For instance, with staggered treatment timing, valid inference conditional on not rejecting pre-trends tests under the null of no pre-trends was put forward as an argument in favor of some of the recently proposed heterogeneity-robust estimators (Borusyak et al., 2024). Actually, our result implies that all heterogeneity-robust estimators lead to exact or conservative inference conditional on not rejecting a pretrends test.
Importantly, our results only hold under the null that the conditions underlying the validity of $\widehat{\beta}$ hold. Thus, they have no bearing on the negative results of Leeb and Pötscher (2005) and Leeb and Pötscher (2006), who show that under local violations of those conditions, pre-testing can distort inference. Under the alternative that the assumptions underlying it are violated, $\widehat{\beta}$ is biased and leads to invalid inference anyways, so the question becomes whether pre-testing, and using $\widehat{\beta}$ when the pre-test is incorrectly not rejected, leads to a larger bias and size distortion than not pre-testing and using $\widehat{\beta}$ anyways. There is probably not a universal answer to that question. In the DID example, Roth (2022) conducts extensive simulations taylored to a large number of published articles in economics. His results suggest on average a small bias aggravation due to pre-testing, though there are also examples where pre-testing reduces the estimator’s bias.
2 The set-up and main result
We are interested in inference on a parameter $\beta_{0}\in\mathbb{R}^{p}$. To this end, we rely on an estimator $\widehat{\beta}$ which is supposed to be consistent for $\beta_{0}$, but only if some conditions are satisfied. We assume to have at our disposal one or several statistical tests of either these conditions themselves, or implications of these conditions. The question we ask is whether inference on $\beta_{0}$ based on $\widehat{\beta}$ is valid when it is performed only conditional on not rejecting any of these tests.
Example 1 (Difference-in-differences)
We assume that some units (the “treated”) receive a binary treatment at a period of time, whereas other units (“the controls”) do not. We are interested in $\beta_{0}$, the average treatment effect on the treated after treatment occurs. $\beta_{0}$ could also correspond to the vector of average treatment effects at different horizons after the treatment takes place. If parallel trends hold, $\beta_{0}$ can be consistently estimated by $\widehat{\beta}$, the usual difference-in-difference estimator. Parallel trends at all periods also imply that $\theta_{0}$, a difference-in-differences estimand on pre-treatment periods, should be equal to 0. We report $\widehat{\beta}$ and confidence intervals on $\beta_{0}$ conditional on not rejecting $\theta_{0}=0$. Note that a common practice is to test separately $\theta_{01}=0$, $\theta_{02}=0$ until $\theta_{0q}=0$. Then, we make inference on $\beta_{0}$ conditional on not rejecting any of these tests.
Example 2 (Instrumental variables)
We seek to estimate $\beta_{0}$, the local average treatment effect of an endogenous binary treatment $D$, using a binary instrumental variable $Z$. Under instrument independence, exclusion restriction and instrument monotonicity, $\beta_{0}$ is equal to the Wald estimand (Imbens and Angrist, 1994). We can test these conditions using for instance Kitagawa’s variance-weighted Kolmogorov-Smirnov test statistic (Kitagawa, 2015). Oftentimes, researchers also perform balancing tests on covariates, as a proxy for testing the instrument independence.
Example 3 (GMM)
We are interested in $\beta_{0}\in\mathbb{R}^{p}$, which satisfies the moment conditions $E[g(U,\beta_{0})]=0$, where $U$ denotes an observed random vector, and $g(u,\beta)\in\mathbb{R}^{q}$, $q\geq p$. If $q>p$, we can test these conditions with an overidentifying test, using either the standard $J$-test or a weighting matrix different from the asymptotically optimal one. In the latter case, the test statistic does not converge to a chi-squared distribution, but we can still perform the test using the bootstrap, for instance. We perform inference in $\beta_{0}$ only when we do not reject the overidentifying test.
Example 4 (Linear and constant effect models with uncounfoundedness)
Consider a continuous treatment $D$ with support $\mathcal{D}$, let $Y(d)$ denote the potential outcome corresponding to $D=d$ and let $Y:=Y(D)$ be the observed outcome. Assume that i) $E(Y(0)|D)=\alpha_{0}$ and ii) $d\mapsto Y(d)=Y(0)+\beta_{0}d$, for two real numbers $(\alpha_{0},\beta_{0})$. Then, $\beta_{0}$ can be estimated by a simple linear regression of $Y$ on $D$, and let $\widehat{\beta}$ denote the corresponding OLS estimator. i) and ii) imply that $E[Y|D]=\alpha_{0}+\beta_{0}D$, a testable implication (de Chaisemartin and D’Haultfœuille, 2024). Define the function $\theta_{0}$ by
$\theta_{0}(d)=E[\mathds{1}\left\{D\leq d\right\}(Y-\alpha_{0}-\beta_{0}D)],$
where $\alpha_{0}=E[Y]-\beta_{0}E[D]$. If $D\perp\!\!\!\perp(Y(d))_{d\in\mathcal{D}}$ and $d\mapsto Y(d)$ is linear, then $d\mapsto E[Y|D=d]$ is linear, which is equivalent to having $\theta_{0}(d)=0$ for all $d\in\mathcal{D}$.
We formalize our problem with the following assumption. Hereafter, we allow for test statistics that are functionals of the empirical measure. We thus use the corresponding standard terminology and notation below; see, e.g., Van der Vaart and Wellner (2023).
@@ 2 The set-up and main result
    We observe a sample $(U_{i})_{i=1,...,n}$ of i.i.d. random vectors with probability distribution $P_{U}$.
@@
    If $P_{U}\in\mathcal{P}_{1}$, with $\mathcal{P}_{0}\subseteq\mathcal{P}_{1}$, we have $Pf:=E[f(U_{1})]=0$ for all $f\in\mathcal{F}$, a Donsker class of real-valued functions. [^1] For the definition of a Donsker class, see, e.g., Van der Vaart and Wellner (2023), p.130.
@@
    We consider $J$ tests of $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
@@
    where $P_{n}$ denotes the empirical measure of $U$, indexed by $\mathcal{F}$. Moreover, $T_{j}$ is convex and satisfies $T_{j}(-P)=T_{j}(P)\geq 0$ and $T_{j}(0)=0$. Finally, $T_{j}$ is continuous with respect to the uniform norm $\|P\|_{\mathcal{F}}:=\sup_{f\in\mathcal{F}}|Pf|$.
  5. 5.
    The critical region of the $j$-th test of $H_{0}$ is $\{T_{j,n}>q_{j,n}\}$, where $q_{j,n}$ is a random variable satisfying $q_{j,n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}q_{j}>0$.
Condition 1 is assumed for simplicity: our results would go through as long as we can establish a uniform central limit theorem. Equation (1) holds under mild regularity restrictions for large classes of estimators (e.g., M- or GMM estimators). Condition 3 states that we can test $H_{0}:\;P_{U}\in\mathcal{P}_{1}$ by checking that some moment conditions hold. Importantly, the class $\mathcal{F}$ may be infinite (and uncountable), in which case we seek to test infinitely many moment restrictions. Conditions 4 and 5 impose some restrictions on the type of tests we are considering. Continuity is typically imposed or checked to derive the weak convergence of $T(P_{n})$ using the continuous mapping theorem. Finally, the last condition imposes the standard condition that the thresholds used in the $J$ tests converge to positive values. The condition $q_{j}>0$ simply rules out that the threshold is at the boundary of the support of the asymptotic distribution of $T_{j,n}$ under $H_{0}$.
We give below three common examples of tests satisfying Conditions 4-5.
Example 5
(F-tests) If $\mathcal{F}$ is finite, =$\mathcal{F}=\{f_{1},...,f_{K}\}$ is Donsker if and only if the functions $f_{1},...,f_{K}$ are measurable and $\max_{k=1,...,K}E[f_{k}(U_{1})^{2}]<\infty$. Let $f=(f_{1},...,f_{K})^{\prime}$ and
$T_{1,n}=n\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})\right)^{\prime}\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})f(U_{i})^{\prime}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}f(U_{i})\right).$
We reject the null hypothesis if $T_{1,n}>q_{1-\alpha}(K)$, the quantile of order $1-\alpha$ of a chi-squared distribution with $K$ degrees of freedom. By Slutsky’s lemma,
$T_{1,n}=\left(n^{1/2}(P_{n}f_{1},...,P_{n}f_{K})\right)E[f(U_{1})f(U_{1})^{\prime}]^{-1}\left(n^{1/2}(P_{n}f_{1},...,P_{n}f_{K})^{\prime}\right)+o_{P}(1).$
Therefore, $T_{1,n}=T_{1}(n^{1/2}P_{n})+o_{P}(1)$, where $T_{1}(Q)=(Qf_{1},...,Qf_{K})E[f(U_{1})f(U_{1})^{\prime}]^{-1}(Qf_{1},...,Qf_{K})^{\prime}$. It is easily seen that $T_{1}$ satisfies Condition 4 of Assumption 1.
Example 6
@@
(Weighted Kolmogorov-Smirnov test statistics) Such test statistics are defined by $T_{1}(P)=\sup_{f\in\mathcal{F}}w(f)|Pf|$ for some Donsker clas $\mathcal{F}$. This includes for instance the test statistic of Kitagawa (2015), in which $\mathcal{F}$ is the set of intervals of the real line. [^2] In his case, $T_{1,n}=\sup_{f\in\mathcal{F}}w_{n}(f)|Pf|$ for a random $w_{n}(\cdot)$, but $w_{n}$ converges uniformly to some deterministic $w$, so (3) holds in his context. If we assume that for all $f\in\mathcal{F}$, $0<\underline{w}\leq w(f)\leq\overline{w}<\infty$, then Condition 4 of Assumption 1 holds. To see that it is continuous, fix $P$ and $\delta>0$ and let $Q$ be such that $\|P-Q\|_{\mathcal{F}}<\delta$. Let $f^{*}$ be such that $w(f^{*})|Pf^{*}|>T_{1}(P)-\delta$. Then
@@
Example 7

### v2

@@ Abstract
Consider a parameter of interest, which can be consistently estimated under some conditions. Suppose also that we can at least partly test these conditions with specification tests. We consider the common practice of conducting inference on the parameter of interest conditional on not rejecting these tests. We show that if the tested conditions hold, conditional inference is valid, though possibly conservative. This holds generally, without imposing any assumption on the asymptotic dependence between the estimator of the parameter of interest and the specification test.
@@ 1 Introduction
Let $\beta_{0}$ denote a parameter of interest. Let $\widehat{\beta}$ denote an estimator of $\beta_{0}$, which is consistent and asymptotically normal if the probability distribution generating the data, $P_{U}$, belongs to some set $\mathcal{P}_{0}$. Finally, assume that the null hypothesis of valid specification, i.e. $P_{U}\in\mathcal{P}_{0}$, is partly testable. In this paper, we consider the common practice of reporting $\widehat{\beta}$ and a confidence interval for $\beta_{0}$ only if a pre-test that $P_{U}\in\mathcal{P}_{0}$ is not rejected. One may wonder if pre-testing distorts inference. This note shows that under the null that $P_{U}\in\mathcal{P}_{0}$, conditional on not rejecting the pre-test the probability that $\beta_{0}$ belongs to its confidence interval (CI) is at least as large as the CI’s nominal level. Therefore, conditional inference is valid, albeit possibly conservative. This result holds under general conditions. In particular, no restriction on the asymptotic dependence between $\widehat{\beta}$ and the pre-tests is needed. The main condition on the pre-tests is that they need to be convex functionals. This holds for finite-dimensional $F$-tests, and for infinite-dimensional tests based on empirical processes such as (weighted) Kolmogorov-Smirnov or Crámer-von Mises tests. For finite-dimensional test statistics, our result is a direct application of the Gaussian correlation inequality (Royen, 2014). For infinite-dimensional ones, we rely on an approximation before applying this inequality.
Our result is relevant for statistical methods using observational data to study cause-and-effect relations. All those methods have to rely on an identifying assumption to recover treated units’ unobserved counterfactual outcome, and researchers typically start by conducting specification tests to establish the credibility of that assumption. For instance, difference-in-difference (DID) estimators rely on a parallel trends assumption, and researchers typically assess the plausibility of that assumption by checking that prior to treatment, treated and control units were experiencing parallel outcome evolutions. Then, they may report the DID estimator and its confidence interval only if the pre-test is not rejected. Our result implies that under the null of parallel trends, inference conditional on not rejecting a pre-trends test is at worst conservative, but can never be liberal, irrespective of the correlation structure between the DID and pre-trends estimators. This complements Proposition 1 in Roth (2022), that shows that under parallel trends, the DID estimator remains conditionally unbiased. This also has implications for some recent discussions in that literature. For instance, with staggered treatment timing, various heterogeneity-robust estimators have been proposed. Borusyak et al. (2024) show, under some assumptions, that under the null of parallel trends their estimator leads to exact inference conditional on not rejecting a pre-trends test, which they view as an advantage over other estimators. Actually, our result implies that under the null of no pre-trends, all heterogeneity-robust estimators lead to exact or conservative inference conditional on not rejecting a pre-trends test. Beyond pre-trends tests in DID studies, our results also apply to continuity tests in regression discontinuity designs, to balancing and specification tests in instrumental variable studies, or to over-identification tests in generalized method of moments estimation, to name a few other examples. [^1] Another example is Angelini et al. (2024), who develop, in the context of proxy-SVARs, a pre-test of the null hypothesis of instrument relevance. This pre-test is asymptotically independent of the main estimators, thus leading to valid conditional inference. Our results show that other pre-tests could also lead to valid albeit possibly conservative inference, even if asymptotic independence does not hold.
Our result shows that under the null of valid specification, pre-tests of valid specification cannot make inference liberal. Under the alternative, namely if $P_{U}\notin\mathcal{P}_{0}$, $\widehat{\beta}$ may be inconsistent, and it may lead to invalid inference anyway. Then, the question becomes whether pre-testing and using $\widehat{\beta}$ when the pre-test is incorrectly not rejected leads to a larger bias and size distortion than not pre-testing and always using $\widehat{\beta}$. In the DID example, Roth (2022) conducts simulations tailored to 12 published articles in economics, assuming that the parallel-trends assumption does not hold due to differential linear trends that have 50% chances of being detected by a pre-trends test. Column 3 of his Table 3 shows that for 10 articles, the bias of $\widehat{\beta}$ conditional on wrongly not-rejecting the pre-trends test is larger than the unconditional bias. However, there are only two articles where the conditional bias is more than 20% larger than the unconditional one, and there are also two articles where the conditional bias is lower. Moreover, the pre-trends test still correctly leads the analyst to not use the DID estimator 50% of the time, which can also lead to a bias reduction if they then use a less-biased estimator instead. Overall, while our result gives a justification for pre-testing under the null, our reading of the simulations in Roth (2022) is that in DID studies, pre-testing is rarely very harmful under the alternative. To our knowledge, no study has investigated whether pre-testing can be harmful under the alternative in other contexts, such as regression discontinuity designs, instrumental variable studies, or generalized method of moments estimation. Doing so is an interesting avenue for future research.
Finally, our paper is related to the literature on inference post model selection (see, e.g. Leeb and Pötscher, 2005, 2006), though the setup is different because in our setting, there may be no estimation taking place when the pre-test is rejected. To understand this, consider a simple model selection set-up with two nested models. Then, the goal of the pre-test is to determine which of the two models will be used to estimate the coefficients on the variables included in the sparser model. The negative results in that literature are under local alternatives, where the true model is the richer one. Then, pre-testing can lead to liberal inference. Instead, under the null that the true model is the sparser one, our result implies that pre-testing cannot make inference liberal, and we also show in Appendix A that this positive result holds uniformly. Our result complements several other positive results in that literature, which have considered cases, such as nested linear models, where the estimator of the sparse model is asymptotically independent of the pre-test, so that conditional inference is exact under the null (see, e.g. Proposition 3.1 in Leeb and Pötscher, 2003). Instead, our result does not need to impose conditions ensuring asymptotic independence, but it only guarantees conservative inference.
2 Set-up and examples
We are interested in a parameter $\beta_{0}\in\mathbb{R}^{p}$. We use an estimator $\widehat{\beta}$, which is consistent and asymptotically normal if the probability distribution generating the data, $P$, belongs to some set $\mathcal{P}_{0}$. Finally, we assume that the null hypothesis of valid specification, i.e. $P\in\mathcal{P}_{0}$, has testable implications. We now review several instances where those conditions are met.
Example 1 (Difference-in-differences, DID)
We have a panel data set with $T>2$ periods. Some units (the “treated”) receive a binary treatment at a period $t_{0}>2$, whereas other units (“the controls”) do not. We are interested in $\beta_{0}$, the average treatment effect on the treated after treatment occurs. $\beta_{0}$ could also be the vector of average treatment effects at $t_{0}$, $t_{0}+1$, …, $T$. In this example, the null hypothesis of valid specification, i.e. $P\in\mathcal{P}_{0}$, is that the average counterfactual outcome evolutions without treatment of treated and control units are the same at all periods. Under that parallel-trends assumption, $\beta_{0}$ can be consistently estimated by $\widehat{\beta}$, the usual DID estimator. This parallel trends assumption also has $t_{0}-2$ testable implications: it implies that $\theta_{0}$, a vector of DID estimands comparing the average outcome evolutions of treated and control units from $t_{0}-1$ to $t_{0}-2$, from $t_{0}-1$ to $t_{0}-3$, …, and from $t_{0}-1$ to $1$, is equal to 0.
Example 2 (Regression discontinuity designs, RDD)
To simplify the discusssion, we focus on sharp RDD. We are interested in the effect of a binary treatment, where treatment occurs if and only if a running variable $X$ is above a threshold, 0 say. The parameter of interest is $\beta_{0}=E[Y(1)-Y(0)|X=0]$, with $Y(d)$ denoting the potential outcome under treatment value $d$. If for all $d$ $x\mapsto E[Y(d)|X=x]$ is continuous at 0, $\beta_{0}:=\lim_{x\downarrow 0}E[Y|X=x]-\lim_{x\uparrow 0}E[Y|X=x]$, which can be estimated by the difference of two nonparametric estimators (Hahn et al., 2001). Two usual specification tests are a test of continuity of the density of $X$ at 0 (McCrary, 2008), and a test that predetermined variables $W$ have continuous means at the threshold. In this example, the null hypothesis of valid specification could for instance be Conditions 1b and 2b in Lee (2008), which imply both the RD identifying assumption ($x\mapsto E[Y(d)|X=x]$ continuous at 0) and the null hypotheses of the two specification tests.
Example 3 (Instrumental variables, IV)
We seek to estimate $\beta_{0}$, the local average treatment effect of an endogenous binary treatment $D$ on an outcome $Y$, using a binary instrumental variable $Z$. In this example, the null hypothesis of valid specification is that the instrument is independent of potential outcomes and treatments, is excluded from the outcome equation, and has a monotonic effect on the treatment. Under those assumptions, $\beta_{0}$ is equal to the Wald estimand (Imbens and Angrist, 1994). Kitagawa (2015) shows that those assumptions have a testable implication:
$P(Y\leq y,D=d|Z=d)-P(Y\leq y,D=1-d|Z=d)\geq 0,$ (1)
for any $y$ in the support of $Y$ and $d\in\{0,1\}$. If $Y$ is continuously distributed, the null of valid specification has infinitely many testable implications, which we can for instance test using the variance-weighted Kolmogorov-Smirnov test statistic proposed in Kitagawa (2015). In IV studies, researchers sometimes also test that some pre-determined covariates $W$ are mean-independent of the instrument, a testable implication of a stronger null where $Z$ is also assumed to be independent of the covariates.
Example 4 (Generalized method of moments)
We are interested in $\beta_{0}\in\mathbb{R}^{p}$, which satisfies the moment conditions $E[g(U,\beta_{0})]=0$, where $U$ denotes an observed random vector, and $g(u,\beta)\in\mathbb{R}^{q}$, $q\geq p$. The moment conditions identify $\beta_{0}$, and if $q>p$, the moment conditions are testable, using a standard J-test (Sargan, 1958; Hansen, 1982).
Example 5 (Linear regressions with a non-binary, uncounfounded treatment)
Let $D$ be a non-binary treatment whose support $\mathcal{D}\subset\mathbb{R}$ includes 00, let $Y(d)$ (resp. $Y$) denote the potential outcome corresponding to $d\in\mathcal{D}$ (resp. the observed outcome), and let $S=(Y-Y(0))/D$ denote units’ potential outcome slope, between a treatment of zero and their actual treatment. We are interested in $\beta_{0}=E[S]$. Researchers sometimes use $\widehat{\beta}$, the coefficient on $D$ in a linear regression of $Y$ on $D$, to estimate $\beta_{0}$. In this example, the null hypothesis of valid specification is that $Y(0)$ and $S$ are mean-independent of $D$. Then, $\widehat{\beta}$ is consistent for $\beta_{0}$. The null hypothesis has the following testable implication:
$E[Y|D]-(\alpha_{0}+\beta_{0}D)=0,$ (2)
where $\alpha_{0}=E[Y]-\beta_{0}E[D]$. In settings where the treatment did not exist before the study period, and where one has access to outcome data $Y_{-1}$ for a prior period, one can placebo test that $Y(0)$ is mean-independent of $D$, for instance by regressing $Y_{-1}$ on $D$. Furthermore, Theorem 1 in de Chaisemartin et al. (2024) shows that under the maintained assumption that $Y(0)$ is mean-independent of $D$, (2) implies that $S$ is mean-independent of $D$. Accordingly, de Chaisemartin et al. (2024) argue that one may report $\widehat{\beta}$ when a placebo test that $Y(0)$ is mean-independent of $D$ and a test of (2) are not rejected. If $D$ is continuously distributed, the null of valid specification has infinitely many testable implications, which we can for instance test using the Kolmogorov-Smirnov or Cramér-von Mises test of linearity of $E[Y|D]$ proposed in Stute (1997) and Stute et al. (1998).
In order to encompass all the examples above, we consider the following two alternative assumptions. In Assumption 2, we allow for test statistics that are functionals of the empirical measure; some of our terminology and notation reflects this.
@@ 2 Set-up and examples
    We observe a sample $(U_{i})_{i=1,...,n}$ of identically distributed random vectors ($U_{i}\in\mathbb{R}^{k}$) with probability distribution $P_{U}(A):=P(U_{1}\in A)$ for all Borel set $A\subset\mathbb{R}^{k}$.
  2. 2.
    If $P_{U}\in\mathcal{P}_{0}$, we have,
    $\left(\widehat{\Sigma}_{\beta}^{-1/2}\left(\widehat{\beta}-\beta_{0}\right),\widehat{\Sigma}_{\theta}^{-1/2}\left(\widehat{\theta}-\theta_{0}\right)\right)\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}\left(0,\begin{pmatrix}I_{p}&\Sigma_{12}\\ \Sigma_{12}^{\prime}&\Sigma_{22}\end{pmatrix}\right),$
    where $I_{p}$ is the identity matrix of size $p$, $\Sigma_{12}$ is a matrix of size $p\times q$ and $\widehat{\Sigma}_{\beta}$ (resp. $\widehat{\Sigma}_{\theta}$ and $\Sigma_{22}$) is symmetric positive of size $p\times p$ (resp. $q\times q$).
  3. 3.
    If $P_{U}\in\mathcal{P}_{0}$, we have $\theta_{0}=0$.
  4. 4.
    We consider $J$ specification tests of $H_{0}:P_{U}\in\mathcal{P}_{0}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
    $T_{j,n}=T_{j}\left(\widehat{\Sigma}_{\theta}^{-1/2}\widehat{\theta}\right)+o_{P}(1).$ (3)
    Moreover, $T_{j}$ is convex on $\mathbb{R}^{q}$ and satisfies $T_{j}(-x)=T_{j}(x)\geq 0$ for all $x\in\mathbb{R}^{q}$ and $T_{j}(0)=0$. The critical region of the $j$-th test is $\{T_{j,n}>q_{j,n}\}$, where $q_{j,n}$ is a random variable satisfying $q_{j,n}\stackrel{{\scriptstyle P}}{{\longrightarrow}}q_{j}>0$.
Assumption 2
  1. 1.
    We observe a sample $(U_{i})_{i=1,...,n}$ of i.i.d. random vectors with probability distribution $P_{U}(A):=P(U_{1}\in A)$ for all Borel set $A$.
@@
    If $P_{U}\in\mathcal{P}_{0}$, we have $Pf:=E[f(U_{1})]=0$ for all $f\in\mathcal{F}$, a Donsker class of real-valued functions. [^2] For the definition of a Donsker class, see, e.g., Van der Vaart and Wellner (2023), p.130.
@@
    We consider $J$ specification tests of $H_{0}:P_{U}\in\mathcal{P}_{0}$ based on the statistics $(T_{1,n},...,T_{J,n})\in\mathbb{R}^{J}$ that satisfy, under $H_{0}$,
@@
[…]
We first discuss the common aspects of Assumptions 1 and 2. First, we do not restrict in either of them the asymptotic dependence of $\widehat{\beta}$ and the test statistics $(T_{1,n},...,T_{J,n})$. Second, we assume in both cases that $\widehat{\beta}$ is consistent for $\beta_{0}$ if $P_{U}\in\mathcal{P}_{0}$, but we only test implications of $P_{U}\in\mathcal{P}_{0}$, rather than $P_{U}\in\mathcal{P}_{0}$ itself. Also, it could be that $\widehat{\beta}$ is actually consistent for $\beta_{0}$ under a weaker condition than $P_{U}\in\mathcal{P}_{0}$. This flexibility is important because the tested conditions are sometimes stronger than necessary to have that $\widehat{\beta}$ is consistent, see for instance Examples 1 and 2. Third, the $J$ different tests in Point 4 of Assumptions 1 and 2 reflect the fact that one often considers different specification tests, rather than a single test. Fourth, in both assumptions the functions $T_{1},...,T_{J}$ in the specification tests are convex and symmetric around zero. This condition is important to obtain our main result. It is typically satisfied in practice: all the examples below meet these requirements. Finally, the condition $q_{j}>0$ at the end of Assumptions 1 and 2 simply ensures that the threshold of the test is not at the boundary of the support of the asymptotic distribution of $T_{j,n}$ under $H_{0}$. Note that we do not require the specification tests to be consistent or that they reach their nominal level asymptotically.
Now, let us discuss the differences between Assumption 1 and Assumption 2 and how they complement each other. On the one hand, Assumption 1 allows for estimators converging at rates lower than $n^{1/2}$, as is the case in the RDD example, and it also allows for non-independent (though identically distributed) variables, thus including the case of stationary time series. Such cases are ruled out in Assumption 2.444However, our result under Assumption 2 would go through as long as we can establish the weak convergence of the empirical process $n^{1/2}(P_{n}-P_{U})$, indexed by $\mathcal{F}$, to a Gaussian process. On the other hand, the specification tests in Assumption 1 are based on a finite-dimensional estimand $\theta_{0}$. This is not the case in Assumption 2, where the class $\mathcal{F}$ in Point 3 may be infinite and even uncountable. The continuity condition in Point 4 of Assumption 2 is important in this potentially infinite-dimensional setup, to apply the continuous mapping theorem.
We now review five examples of specification tests. The first two fall under Assumption 1, while the last three fall under Assumption 2.
Specification Test 1 (DID)
In the DID example, for all $t\in\{1,...,t_{0}-2\}$ $\widehat{\theta}_{t}$ is a DID estimator $t$ periods prior to treatment, and $\widehat{\theta}=(\widehat{\theta}_{1},...,\widehat{\theta}_{t_{0}-2})$. Then, a so-called F-test for pre-trends corresponds to the case where $J=1$, $\widehat{\Sigma}_{\theta}$ is a consistent estimator of the asymptotic variance of $\widehat{\theta}$, $T_{1}(x)=x^{\prime}x$, $T_{1,n}=T_{1}(\widehat{\Sigma}_{\theta}^{-1/2}\widehat{\theta})$, and $q_{1,n}$ is the quantile of a chi-squared distribution with $q$ degrees of freedom. Similarly, the so-called “Sup-t” test proposed by Montiel Olea and Plagborg-Møller (2019) corresponds to the case where $J=1$, $\widehat{\Sigma}_{\theta}$ is as before, for all $(x_{1},...,x_{t_{0}-2})\in\mathbb{R}^{t_{0}-2}$ $T_{1}(x_{1},...,x_{t_{0}-2})=\max(|x_{1}|,...,|x_{t_{0}-2}|)$, $T_{1,n}=T_{1}\left(\widehat{\theta}_{1}/\widehat{\Sigma}_{\theta,1,1},...,\widehat{\theta}_{t_{0}-2}/\widehat{\Sigma}_{\theta,t_{0}-2,t_{0}-2}\right)$ with $\widehat{\Sigma}_{\theta,j,j}$ denoting the $j$th diagonal element of $\widehat{\Sigma}_{\theta}$, and $q_{1,n}$ is the quantile of the max of $t_{0}-2$ normally distributed variables with mean 0 and variance $\widehat{\Sigma}_{\theta}$.
Specification Test 2 (RDD)
To test the continuity of the density of $X$ at 0, one uses $\widehat{\theta}_{1}:=\lim_{x\downarrow 0}\widehat{f}_{X}(x)-\lim_{x\uparrow 0}\widehat{f}_{X}(z)$, with $\widehat{f}_{X}$ a nonparametric (e.g., kernel) estimator of $f_{X}$. To test that a predetermined scalar variable $W$ has a continuous mean at the threshold, one uses $\widehat{\theta}_{2}:=\lim_{x\downarrow 0}\widehat{E}[W|X=x]-\lim_{x\uparrow 0}\widehat{E}[W|X=x]$, with $\widehat{E}[W|X=x]$ a nonparametric (e.g., local linear) estimator of $E[W|X=x]$. Then, $\widehat{\theta}=(\widehat{\theta}_{1},\widehat{\theta}_{2})^{\prime}$. One may conduct the two specification tests separately, without adjusting for multiple testing. Then, $J=2$, $\widehat{\Sigma}_{\theta}$ is diagonal with $j$-th diagonal term equal to a consistent estimator of the asymptotic variance of $\widehat{\theta}_{j}$, $T_{j}(x)=x_{j}^{\prime}x_{j}$, $T_{j,n}=T_{j}(\widehat{\Sigma}_{\theta}^{-1/2}\widehat{\theta})$ and $q_{j,n}$ is the quantile of a chi-squared distribution with one degree of freedom.
Specification Test 3 (Weighted Kolmogorov-Smirnov test statistics)
@@
Such test statistics correspond to $T_{1}(P)=\sup_{f\in\mathcal{F}}w(f)|Pf|$ for some Donsker clas $\mathcal{F}$. This includes for instance the test statistic of Kitagawa (2015), used in Example 3, and where $\mathcal{F}$ is the set of intervals of the real line. [^5] In his case, $T_{1,n}=\sup_{f\in\mathcal{F}}w_{n}(f)|Pf|$ for a random $w_{n}(\cdot)$, but $w_{n}$ converges uniformly to some deterministic $w$, so (5) holds in his context. If we assume that for all $f\in\mathcal{F}$, $0<\underline{w}\leq w(f)\leq\overline{w}<\infty$, then $T_{1}$ satisfies the restrictions in Point 4 of Assumption 2. To see that it is continuous, fix $P$ and $\delta>0$ and let $Q$ be such that $\|P-Q\|_{\mathcal{F}}<\delta$. Let $f^{*}$ be such that $w(f^{*})|Pf^{*}|>T_{1}(P)-\delta$. Then,
@@
By considering $f^{**}$ be such that $w(f^{**})|Qf^{**}|>T_{1}(Q)-\delta$, we obtain the same inequality for $T_{1}(Q)-T_{1}(P)$. Hence, $|T_{1}(P)-T_{1}(Q)|<(\overline{w}+1)\;\delta$, which implies that $T_{1}$ is continuous.
Specification Test 4 (Generalized Crámer-von Mises statistic)